Org Pvt. Ltd. is considering two mutually exclusive capital investments. The project’s

expected net cash flows are as follows

Expected Cash Flows

Year Project A Project B

0 -400 -575

1 95 150

2 110 200

3 118 250

4 125 275

5 140 230

6 150 180


If you were told that each project’s cost of capital was 10%, which project should be selected using the NPV criteria?

b.What is each project’s IRR?

c.What is the regular payback period for these two projects?

d.What is the profitability index for each project if the cost of capital is 12%?


1
Expert's answer
2021-05-14T12:03:46-0400

A. Project A NPV = 1138- 113.8=1024.2

Project B NPV= 1860- 186= 1674

Project B should be selected.

B. Project A IRR= 184.5%

Prokect B IRR= 233.5%

C. Project A payback period= 2.8 years

Project B payback period= 3.2 years

D. Project A= 2.5%

Project B= 2.8%